Taylor is 18 years older than Jeniffer. In four years Taylor will be 3 times as old as Jeniffer. What are there ages now ?
Answer by TimothyLamb(4379)

---
x = y + 18
x + 4 = 3(y + 4)
---
put the system of linear equations into standard form
---
x - y = 18
x + 4 = 3y + 12
---
x - y = 18
x - 3y = 8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= Taylor = 23
y= Jeniffer = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php